OpenStudy (anonymous):

The length of a rectangle is 8 inches greater than its width. Find the width of the rectangle if its area is 105 square inches.@noemy_bieber

OpenStudy (anonymous):

 Area of a rectangle = length * width width = x length = x + 8 therefore Area of the rectangle = (x) ( x+8) 105 = (x)(x+8)